OpenStudy (anonymous):

A car has a mass of 1.20 × 103 kilograms and a momentum of 2.00 × 104 kilogram meters/second. What is the velocity of the car?

OpenStudy (anonymous):

Formula for momentum: P = 1/2(m)(v^2) Sub in givens (2.00 x 10^4) = 1/2 x (1.20 x 10^3) (v ^2) rearrange and you get \[v = \sqrt{ 2*(20000 / 1200)}\]

OpenStudy (anonymous):

Alrighty Ms. SmartyPants, so my answer is 5.77 m/s? Or am I wrong? @biggi3lue

OpenStudy (anonymous):

Perfect! Well done. Now, with this knowledge you can estimate how fast something will be moving after a car crash! The only thing you need to account for is the cars crumple zone which absorbs the momentum energy. But for something like a tennis ball, if it were to somehow magically absorb all the momentum, it would travel something like 1000kilometres/second! WOW

OpenStudy (anonymous):

Dang!(: Thank you! @biggi3lue